9x-3y=24
7x-3y=20
using the elimination method

Answer by funmath(2933)

9x-3y=24
7x-3y=20
using the elimination method
Multiply the second equation by -1 and add it to the first equation and the y's will be eliminated so that you can solve for x:
-1(7x-3y)=-1(20)
-7x+3y=-20
:
9x-3y=24
-7x+3y=-20
----------
2x+0=4
2x=4
2x/2=4/2
x=2
Substitute x=2 into either equation and solve for y:
9(2)-3y=24
18-3y=24
18-18-3y=24-18
-3y=6
-3y/-3=6/-3
y=-2
The solution (x,y)=(2,-2)
Happy Calculating!!!